1.   Subtract the binary numbers 11102 from 101012

1
Expert's answer
2021-08-16T16:35:51-0400

Let us subtract the binary numbers 1110from 101012:


"10101_2\\
\ \ 1110_2\\
---\\
\ \ \ \ 111_2"


Let us verify using decimal system:


"10101_2=1\cdot 2^4+0\cdot 2^3+1\cdot 2^2+0\cdot 2+1=16+4+1=21_{10} \\
1110_2=1\cdot 2^3+1\cdot 2^2+1\cdot 2+0=8+4+2=14_{10}\\
111_2=1\cdot 2^2+1\cdot 2+1=4+2+1=7_{10}"


And "14+7=21"


Answer: 1112